# Laude: LLM-Assisted Unit Test Generation and Debugging of Hardware DEsigns

### v1

@@ Abstract
Unit tests are critical in the hardware design lifecycle to ensure that component design modules are functionally correct and conform to the specification before they are integrated at the system level. Thus developing unit tests targeting various design features requires deep understanding of the design functionality and creativity. When one or more unit tests expose a design failure, the debugging engineer needs to diagnose, localize, and debug the failure to ensure design correctness, which is often a painstaking and intense process. In this work, we introduce Laude, a unified unit-test generation and debugging framework for hardware designs that cross-pollinates the semantic understanding of the design source code with the Chain-of-Thought (CoT) reasoning capabilities of foundational Large-Language Models (LLMs). Laude integrates prompt engineering and design execution information to enhance its unit test generation accuracy and code debuggability. We apply Laude with closed- and open-source LLMs to a large corpus of buggy hardware design codes derived from the VerilogEval dataset, where generated unit tests detected bugs in up to 100% and 93% of combinational and sequential designs and debugged up to 93% and 84% of combinational and sequential designs, respectively.
@@
Deeksha Nandal1thanks: DN and RR equally contributed in this work., Riccardo Revalor1, Soham Dan2thanks: SD and DP jointly supervised this work., Debjit Pal [^1] footnotemark: 1 1Dept. of Electrical and Computer Engineering, University of Illinois Chicago, Chicago IL 60607, 2Microsoft, Correspondence: dpal2@uic.edu
@@ 12 State <= g1;
Consider 1-bit full adder (FA) of Figure˜1(a), the FSM of a 2-port arbiter (Arb) of Figure˜1(b), and a buggy Arb of Figure˜1(c). For FA, outputs (s, cout) depend only on inputs (a, b, c) much like a software function. However, for the Arb, the outputs (g1, g2) depends on inputs (r1, r2, g1, g2), State, clock (clk), and an asynchronous reset (rst). Lines 22 – 23 contain the bug that causes Arb to diverge from the desired behavior of the FSM. However, that buggy code will be executed only when State = 0. One would require a unit test to drive the Arb to State = 0 to sensitize the bug requiring creativity and deep undestanding of the Arb behavior captured by the FSM.
Recent research has explored using LLMs for various hardware design and verification tasks to improve syntactic correctness and functional coherence. It remains underexplored whether LLMs can reason and generate test inputs that can trigger and propagate the bugs to observable outputs. This is particularly challenging since it requires reasoning across clock cycles, interactions in FSMs, and is heavily dependent on an implementation (c.f.Figure˜1). Additionally, a unit test, upon failure, should produce a trace that is sufficiently diverge and contain distinct failure signature for successful debugging. However, there is an acute lack of research that explores the reasoning abilities of LLMs to debug design by developing unit tests.
@@ 2 Background and Preliminaries
VerilogEval Dataset. VerilogEval is a hardware design benchmarking suite targeted toward evaluating LLM’s performance in the context of hardware design and verification Liu et al. (2023). VerilogEval consists of 156 design problems from the Verilog instructional website HDLBits segregated into a set of 82 combinational designs and 74 sequential designs with complex finite state machines (FSMs). Each design problem consists of a reference SV code implementation and a reference testbench to simulate the correct design behavior.
@@ 3 Laude Methodology
Problem Statement. Given a natural language description $\mathcal{L}$ of a hardware design $\mathcal{D}$ and an oracle $\mathbb{O}$, we generate a unit test $\mathcal{U}$ such that when $\mathcal{U}$ applied to $\mathcal{D}$, it produces an output that is consistent with $\mathbb{O}$’s output on $\mathcal{U}$ (i.e., $\mathcal{T}_{p}$). Our problem has the following foundations – (i) we consider a parameterized unit test $\mbox{{$\mathcal{U}$}}(i_{1},i_{2},\ldots,i_{m})$ where we treat each design input $i_{k}\in\mathcal{I}$ as a parameter and assign a set of binary values to each input, and (ii) we establish functional correctness of the design by simulating it with a set of $\mbox{{$\mathcal{U}$}}(i_{1},i_{2},\ldots,i_{m})$ and ensuring the output matches the expected output. For brevity, we will denote $\mbox{{$\mathcal{U}$}}(i_{1},i_{2},\ldots,i_{m})$ as $\mathcal{U}$.
@@ 3.1 Unit Test Generation Technique
We consider a unit test $\mathcal{U}$ to be valid for design $\mathcal{D}$ if $\mbox{{$\mathcal{D}_{\mathcal{C}}$}}(\mbox{{$\mathcal{U}$}})\mathrel{\leadsto}\mbox{{$\mathcal{T}_{p}$}}$. We plan to develop a unit test generator $\mathbf{G}_{\Theta}$ with hyperparameters $\Theta$ to generate a valid unit test $\mathcal{U}$ such that $\mbox{{$\mathbf{G}_{\Theta}$}}(\mathcal{L})\mapsto\mbox{{$\mathcal{U}$}}$ without any designer intervention. However, given that a hardware design has a specific set of inputs and output signals (e.g., request, data, acknowledge, etc.) many of which can be of multibits (e.g., a data bus of width 8 bits), it is important that we provide the design input/output description (signature) $\mathcal{S}_{\mathcal{D}}$ as an input to the generator $\mbox{{$\mathbf{G}_{\Theta}$}}(\mathcal{L},\mbox{{$\mathcal{S}_{\mathcal{D}}$}})$. Such signatures can be easily obtained from the design source code via lightweight source code parsing Mike Popolosk (2024). Since the key usage of a unit test is to establish the functional correctness of a specific HDL implementation of a hardware design, we include the (potentially) buggy implementation $\mathcal{D}_{\mathcal{B}}$ in the generator input $\mbox{{$\mathbf{G}_{\Theta}$}}(\mathcal{L},\mbox{{$\mathcal{S}_{\mathcal{D}}$}},\mbox{{$\mathcal{D}_{\mathcal{B}}$}})$. The intuition here is that by exposing the generator $\mathbf{G}_{\Theta}$ to $\mathcal{D}_{\mathcal{B}}$, it can semantically analyze the source code and identify a set of binary input values that will produce a failure trace exposing the deviation from the design intent as expressed in the description $\mathcal{L}$, i.e., $\mbox{{$\mathcal{D}_{\mathcal{B}}$}}(\mbox{{$\mathcal{U}$}})\mathrel{\leadsto}\mbox{{$\mathcal{T}_{f}$}}$. We use various closed and open source foundational LLMs for $\mathbf{G}_{\Theta}$ (c.f.Section˜4). A key objective of the unit test generation technique is that the unit test should produce a significant divergent trace $\mathcal{T}_{f}$ compared to the output of the $\mathbb{O}$ when applied to the buggy code $\mathcal{D}_{\mathcal{B}}$. In our analysis, we evaluate and compare $\mbox{{$\mathbf{G}_{\Theta}$}}(\mathcal{L})$ (NL), $\mbox{{$\mathbf{G}_{\Theta}$}}(\mathcal{L},\mbox{{$\mathcal{S}_{\mathcal{D}}$}})$ (NLS), and $\mbox{{$\mathbf{G}_{\Theta}$}}(\mathcal{L},\mbox{{$\mathcal{S}_{\mathcal{D}}$}},\mbox{{$\mathcal{D}_{\mathcal{B}}$}})$ (NLSC) as generator techniques.
@@ 4 Experimental Setup
Figure 6: Divergence Rate (DR) distribution for Gemini-2.5 Pro. The $10\times 10$ box glyph heatmap with sparklines utilize equally spaced 5-bin scale: Bin 1 (0%–20%) through Bin 5 (80%–100%) from left to right. The darkest bar in the sparkline indicates the bin containing the median value, which also determines the cell’s background color. NL shows weakest performance. NLS and NLSC achieve the best results, primarily concentrated at Bin 5 (80%–100%) regardless of the testbench source.
@@
Hyperparameters. For all test generator and debugging techniques, we maintained consistent hyperparameters ($\Theta$) to ensure fair comparison – temperature set to 0.8 to balance creativity with determinism. The output token generation limits were set based on the deployment infrastructure. For open-source models, we used a dual strategy to optimize the throughput, consisting of 2,048 for NLSC (to accommodate code-specific reasoning), and 512 for NLS and NL. For closed-source models, we used the API’s maximum token limit, delegating resource management to the provider’s infrastructure. We restricted the input context window to $16,384$ tokens for all open-source models. Although, this number is well below the model’s theoretical maximum taken, we found it to be sufficient to contain the largest design and testbench codes in our datasets with reasonable Key-Value cache memory footprint with efficient execution without Out-of-Memory errors. We set the maximum GPU memory utilization to 0.95 for vLLM.
Figure 7: Divergence Rate (DR) Distribution for DeepSeek R1 using $10\times 10$ box glyph heatmap with 5-bin sparklines as Figure˜6. NL is predominantly concentrated in Bin 1 (0% – 20%). The NLS shows mixed results with high variance. NLSC demonstrates better performance, particularly on targets BC08 – BC10.
@@
Attack Rate (AR). Attack Rate quantifies the bug detectability of a unit test $\mathcal{U}$ when applied on a $\mathcal{D}_{\mathcal{B}}$ and is computed as the fraction of problems for which a $\mathcal{T}_{f}$ is produced. The AR for $i^{th}$ problem is computed as $AR_{i}=\mathbf{1}$ if $\mathcal{T}_{f}$, otherwise $\mathbf{0}$. The overall attack rate is computed as $AR=\frac{\sum_{i=1}^{N}AR_{i}}{N}$, where N is the total number of problems. A higher AR indicates effectiveness of unit tests in bug detection.
Divergent Rate (DR). Divergent Rate quantifies how different/divergent a failure trace $\mathcal{T}_{f}$ is as compared to a passing trace $\mathcal{T}_{p}$ when the same unit test $\mathcal{U}$ is applied on a likely buggy design $\mathcal{D}_{\mathcal{B}}$ and an oracle $\mathbb{O}$, repsectively. DR is measured as $DR=\frac{\sum_{i=1}^{n}|\mbox{{$\mathcal{T}_{p}$}}[i]\neq\mbox{{$\mathcal{T}_{f}$}}[i]|}{n}$, where $\mbox{{$\mathcal{U}$}}||=n$. A higher DR value indicates significant amount of mismatch between $\mathcal{T}_{p}$ and $\mathcal{T}_{f}$, yielding an informative failing trace for debugging. Note while AR measures bug detectability, DR measures bug debuggability.
Divergent Attack (DA). Divergent Attack quantifies the bug detectability and debuggability of a unit test $\mathcal{U}$ in a unified way. Intuitively, a test should be capable of detecting a broad spectrum of bugs (i.e., sensitive to bugs) and generate unique failure signature for a bug to aid debugging (i.e., specific to the effects of the bug). Consequently, we compute DA as $DA=DR\cap AR$, i.e., a unit test $\mathcal{U}$ will have higher DA if it has both high AR and DR.
@@ 5.1 Bug Sensitivity Analysis of Unit Tests
The results show different behaviors across the two LLMs. For Gemini-2.5 Pro, NL is highly effective, consistently achieving AR values falling in the fifth bin ($80\%$). This performance remains stable in NLS and NLSC, with the exception of NLS, where the AR for Source BC09 against Target BC07 drops significantly to $20\%$. DeepSeek R1 is ineffective with NL; AR distributions remain largely in the lowest quartile ($20\%$), except for specific targets (BC03 and BC08) which show higher susceptibility. However, the NLS and NLSC configurations drastically improve DeepSeek R1’s performance, raising the AR to levels that are competitive with Gemini Pro ($80\%$) across all targets. This experiment demonstrates that contextual grounding is a critical requirement for DeepSeek R1 to overcome generation volatility and achieve competitive bug sensitivity.
Figure 8: Divergent Attack (DA) Distribution for Gemini-2.5 Pro using $10\times 10$ box glyph heatmap with 5-bin sparklines. NLS and NLSC achieve the highest results, especially for BC08 – BC10. For BC01 – BC07, we observe higher variance, as DA falls into multiple distinct bins.
@@ 5.2 Bug Specificity Analysis of Unit Tests
For NL, Gemini-2.5 Pro performs the worst characterized by distributions spreading more evenly across intermediate bins. However, Gemini-2.5 Pro demonstrates consistency across the NLS and NLSC, achieving up to 100% (average 90%) DR for BC08 – BC10. However, NLS shows a specific vulnerability for BC02 and BC07, exhibiting a performance drop. This shows that while buggy code generally aids LLM’s unit test generation, specific bug types may act as adversarial examples, degrading the model’s ability to generate unit tests for divergent failure trace.
DeepSeek R1 shows a high sensitivity to the unit test generation strategy. The NL is concentrated in Bin 1 ($0\%-20\%$), indicating a near-total inability to generate divergent traces without context. The NLS shows high variance, struggling to achieve the high-performance of Gemini-2.5. The NLSC shows an improved ability to generate divergent traces, particularly for BC08 – BC10. This experiment shows that the DeepSeek R1 heavily rely on few-shot examples to reason and achieve competitive performance.

### v2

@@ Abstract
Unit tests are critical in the hardware design lifecycle to ensure that component design modules are functionally correct and conform to the specification before they are integrated at the system level. Thus developing unit tests targeting various design features requires deep understanding of the design functionality and creativity. When one or more unit tests expose a design failure, the debugging engineer needs to diagnose, localize, and debug the failure to ensure design correctness, which is often a painstaking and intense process. In this work, we introduce Laude, a unified unit-test generation and debugging framework for hardware designs that cross-pollinates the semantic understanding of the design source code with the Chain-of-Thought (CoT) reasoning capabilities of foundational Large-Language Models (LLMs). Laude integrates prompt engineering and design execution information to enhance its unit test generation accuracy and code debuggability. We apply Laude with closed and open-source LLMs to a large corpus of buggy hardware design codes derived from the VerilogEval dataset, where generated unit tests detected bugs in up to 100% and 93% of combinational and sequential designs and debugged up to 93% and 84% of combinational and sequential designs, respectively.
@@
Deeksha Nandal1thanks: Equal contribution by DN and RR., Riccardo Revalor1, Soham Dan2thanks: SD and DP jointly supervised this work., Debjit Pal [^1] footnotemark: 1 1Dept. of Electrical and Computer Engineering, University of Illinois Chicago, Chicago IL 60607, 2Microsoft, Correspondence: dpal2@uic.edu
@@ 12 State <= g1;
Consider 1-bit full adder (FA) of Figure˜1(a), the FSM of a 2-port arbiter (Arb) of Figure˜1(b), and a buggy Arb of Figure˜1(c). For FA, outputs (s, cout) depend only on inputs (a, b, c), much like a software function. However, for the Arb, the outputs (g1, g2) depends on inputs (r1, r2, g1, g2), State, clock (clk), and an asynchronous reset (rst). Lines 22 – 23 contain the bug that causes Arb to diverge from the desired behavior of the FSM. However, that buggy code will be executed only when State = 0. One would require a unit test to drive the Arb to State = 0 to sensitize the bug, requiring creativity and deep understanding of the Arb behavior captured by the FSM.
Recent research has explored using LLMs for various hardware design and verification tasks to improve syntactic correctness and functional coherence. It remains underexplored whether LLMs can reason and generate test inputs that can trigger and propagate the bugs to observable outputs. This is challenging since it requires reasoning across clock cycles, interactions in FSMs, and is heavily dependent on an implementation (c.f.Figure˜1). Additionally, a unit test, upon failure, should produce a trace that is sufficiently divergent and contain a distinct failure signature for successful debugging. However, there is an acute lack of research that explores the reasoning abilities of LLMs to debug design by developing unit tests.
@@ 2 Background and Preliminaries
VerilogEval Dataset. VerilogEval is a hardware design benchmarking suite targeted toward evaluating LLM’s performance in the context of hardware design and verification Liu et al. (2023). VerilogEval consists of 156 design problems from a Verilog instructional website HDLBits (2025) segregated into a set of 82 combinational designs and 74 sequential designs with complex finite state machines (FSMs). Each design problem consists of a reference SV code implementation and a reference testbench to simulate the correct design behavior.
@@ 3 Laude Methodology
Problem Statement. Given a natural language description $\mathcal{L}$ of a hardware design $\mathcal{D}$ and an oracle $\mathbb{O}$, we generate a unit test $\mathcal{U}$ such that when $\mathcal{U}$ applied to $\mathcal{D}$, it produces an output that is consistent with $\mathbb{O}$’s output on $\mathcal{U}$ (i.e., $\mathcal{T}_{p}$). Our problem has the following assumptions – (i) we consider a parameterized unit test $\mbox{{$\mathcal{U}$}}(i_{1},i_{2},\ldots,i_{m})$ where we treat each design input $i_{k}\in\mathcal{I}$ as a parameter and assign a set of binary values to each input, and (ii) we establish functional correctness of the design by simulating it with a set of $\mbox{{$\mathcal{U}$}}(i_{1},i_{2},\ldots,i_{m})$ and ensuring the output matches the expected output. For brevity, we will denote $\mbox{{$\mathcal{U}$}}(i_{1},i_{2},\ldots,i_{m})$ as $\mathcal{U}$.
@@ 3.1 Unit Test Generation Technique
We consider a unit test $\mathcal{U}$ to be valid for design $\mathcal{D}$ if $\mbox{{$\mathcal{D}_{\mathcal{C}}$}}(\mbox{{$\mathcal{U}$}})\mathrel{\leadsto}\mbox{{$\mathcal{T}_{p}$}}$. We plan to develop a unit test generator $\mathbf{G}_{\Theta}$ with hyperparameters $\Theta$ to generate a valid unit test $\mathcal{U}$ such that $\mbox{{$\mathbf{G}_{\Theta}$}}(\mathcal{L})\mapsto\mbox{{$\mathcal{U}$}}$ without any designer intervention. However, given that a hardware design has a specific set of inputs and output signals (e.g., request, data, acknowledge, etc.) many of which can be of multibits (e.g., a data bus of width 8 bits), it is important that we provide the design input/output description (signature) $\mathcal{S}_{\mathcal{D}}$ as an input to the generator $\mbox{{$\mathbf{G}_{\Theta}$}}(\mathcal{L},\mbox{{$\mathcal{S}_{\mathcal{D}}$}})$. Such signatures can be easily obtained from the design source code via lightweight source code parsing Mike Popolosk (2024). Since the key usage of a unit test is to establish the functional correctness of a specific HDL implementation of a hardware design, we include the (potentially) buggy implementation $\mathcal{D}_{\mathcal{B}}$ in the generator input $\mbox{{$\mathbf{G}_{\Theta}$}}(\mathcal{L},\mbox{{$\mathcal{S}_{\mathcal{D}}$}},\mbox{{$\mathcal{D}_{\mathcal{B}}$}})$. The intuition here is that by exposing the generator $\mathbf{G}_{\Theta}$ to $\mathcal{D}_{\mathcal{B}}$, it can semantically analyze the source code and identify a set of binary input values that will produce a failure trace exposing the deviation from the design intent as expressed in the description $\mathcal{L}$, i.e., $\mbox{{$\mathcal{D}_{\mathcal{B}}$}}(\mbox{{$\mathcal{U}$}})\mathrel{\leadsto}\mbox{{$\mathcal{T}_{f}$}}$. We use various closed and open source foundational LLMs for $\mathbf{G}_{\Theta}$ (c.f.Section˜4). A key objective of the unit test generation technique is that the unit test should produce a significant divergent trace $\mathcal{T}_{f}$ compared to the output of the $\mathbb{O}$ when applied to the buggy code $\mathcal{D}_{\mathcal{B}}$. In our analysis, we evaluate and compare $\mbox{{$\mathbf{G}_{\Theta}$}}(\mathcal{L},\mbox{{$\mathcal{S}_{\mathcal{D}}$}})$ (NLS) and $\mbox{{$\mathbf{G}_{\Theta}$}}(\mathcal{L},\mbox{{$\mathcal{S}_{\mathcal{D}}$}},\mbox{{$\mathcal{D}_{\mathcal{B}}$}})$ (NLSC) as generator techniques. Our initial experiments showed that NL often generates unit tests that are syntactically and/or semantically incorrect and cannot be simulated. In future work, we will further investigate NL and will compare it with NLS and NLSC.
@@ 4 Experimental Setup
Figure 6: Divergence Rate (DR) distribution for Gemini-2.5 Pro. The $10\times 10$ box glyph heatmap with sparklines utilize equally spaced 5-bin scale: Bin 1 (0%–20%) through Bin 5 (80%–100%) from left to right. The darkest bar in the sparkline indicates the bin containing the median value, which also determines the cell’s background color. NLS@0 shows weakest performance. NLSC@0 and NLSC@5 achieve the best results, primarily concentrated at Bin 5 (80%–100%) regardless of the testbench source.
@@
Hyperparameters. For all test generator and debugging techniques, we maintained consistent hyperparameters ($\Theta$) to ensure fair comparison – temperature set to 0.8 to balance creativity with determinism. The output token generation limits were set based on the deployment infrastructure. For open-source models, we used a dual strategy to optimize the throughput, consisting of 2,048 for NLSC (to accommodate code-specific reasoning), and 512 for NLS and NL. For closed-source models, we used the API’s maximum token limit, delegating resource management to the provider’s infrastructure. We restricted the input context window to $16,384$ tokens for all open-source models. Although, this number is well below the model’s theoretical maximum taken, we found it to be sufficient to contain the largest design and testbench codes in our datasets with reasonable Key-Value (KV) cache memory footprint with efficient execution without Out-of-Memory errors. We set the maximum GPU memory utilization to 0.95 for vLLM. The remaining 0.05 of VRAM memory acts as a swap space for temporary tensors and KV pairs.
Figure 7: Divergence Rate (DR) Distribution for DeepSeek R1 using $10\times 10$ box glyph heatmap with 5-bin sparklines as Figure˜6. NLS@0 is predominantly concentrated in Bin 1 (0% – 20%). The NLSC@0 shows mixed results with high variance. NLSC@5 demonstrates better performance, particularly on targets BC08 – BC10.
@@
Attack Rate (AR). Attack Rate quantifies the bug detectability of a unit test $\mathcal{U}$ when applied on a $\mathcal{D}_{\mathcal{B}}$ and is computed as the fraction of problems for which a $\mathcal{T}_{f}$ is produced. The AR for $i^{th}$ problem is computed as $AR_{i}=\mathbf{1}$ if $\mathcal{T}_{f}$, otherwise $\mathbf{0}$. The overall attack rate is computed as $AR=\frac{\sum_{i=1}^{N}AR_{i}}{N}$, where N is the total number of problems. A higher AR indicates the effectiveness of unit tests in bug detection.
Divergent Rate (DR). Divergent Rate quantifies how different/divergent a failure trace $\mathcal{T}_{f}$ is as compared to a passing trace $\mathcal{T}_{p}$ when the same unit test $\mathcal{U}$ is applied on a likely buggy design $\mathcal{D}_{\mathcal{B}}$ and an oracle $\mathbb{O}$, repsectively. DR is measured as $DR=\frac{\sum_{i=1}^{n}|\mbox{{$\mathcal{T}_{p}$}}[i]\neq\mbox{{$\mathcal{T}_{f}$}}[i]|}{n}$, where $\mbox{{$\mathcal{U}$}}||=n$. A higher DR value indicates a significant amount of mismatch between $\mathcal{T}_{p}$ and $\mathcal{T}_{f}$, yielding an informative failing trace for debugging. Note while AR measures bug detectability, DR measures bug debuggability.
Divergent Attack (DA). Divergent Attack quantifies the bug detectability and debuggability of a unit test $\mathcal{U}$ in a unified way. Intuitively, a test should be capable of detecting a broad spectrum of bugs (i.e., sensitive to bugs) and generate a unique failure signature for a bug to aid debugging (i.e., specific to the effects of the bug). Consequently, we compute DA as $DA=DR\cap AR$, i.e., a unit test $\mathcal{U}$ will have higher DA if it has both high AR and DR.
Experimental Categories: We primarily compare 0-shot of NLS (denoted as NLS@0) and 0-shot and 5-shot of NLSC, denoted as NLSC@0 and NLSC@5, respectively.
@@ 5.1 Bug Sensitivity Analysis of Unit Tests
The results show different behaviors across the two LLMs. For Gemini-2.5 Pro, NLS@0 is highly effective, consistently achieving AR values falling in the fifth bin ($80\%$). This performance remains stable in NLSC@0 and NLSC@5, with the exception of NLSC@0, where the AR for Source BC09 against Target BC07 drops significantly to $20\%$. DeepSeek R1 is ineffective with NLS@0; AR distributions remain largely in the lowest quartile ($20\%$), except for specific targets (BC03 and BC08) which show higher susceptibility. However, the NLSC@0 and NLSC@5 configurations drastically improve DeepSeek R1’s performance, raising the AR to levels that are competitive with Gemini Pro ($80\%$) across all targets. This experiment demonstrates that contextual grounding is a critical requirement for DeepSeek R1 to overcome generation volatility and achieve competitive bug sensitivity.
Figure 8: Divergent Attack (DA) Distribution for Gemini-2.5 Pro using $10\times 10$ box glyph heatmap with 5-bin sparklines. NLSC@0 and NLSC@5 achieve the highest results, especially for BC08 – BC10. For BC01 – BC07, we observe higher variance, as DA falls into multiple distinct bins.
@@ 5.2 Bug Specificity Analysis of Unit Tests
For NLS@0, Gemini-2.5 Pro performs the worst characterized by distributions spreading more evenly across intermediate bins. However, Gemini-2.5 Pro demonstrates consistency across the NLSC@0 and NLSC@5, achieving up to 100% (average 90%) DR for BC08 – BC10. However, NLSC@0 shows a specific vulnerability for BC02 and BC07, exhibiting a performance drop. This shows that while buggy code generally aids LLM’s unit test generation, specific bug types may act as adversarial examples, degrading the model’s ability to generate unit tests for divergent failure trace.
DeepSeek R1 shows a high sensitivity to the unit test generation strategy. The NLS@0 is concentrated in Bin 1 ($0\%-20\%$), indicating a near-total inability to generate divergent traces without context. The NLSC@0 shows high variance, struggling to achieve the high-performance of Gemini-2.5. The NLSC@5 shows an improved ability to generate divergent traces, particularly for BC08 – BC10. This experiment shows that the DeepSeek R1 heavily rely on few-shot examples to reason and achieve competitive performance.